# Swin2-MoSE: A New Single Image Super-Resolution Model for Remote Sensing

### v1

@@ Abstract
Due to the limitations of current optical and sensor technologies and the high cost of updating them, the spectral and spatial resolution of satellites may not always meet desired requirements. For these reasons, Remote-Sensing Single-Image Super-Resolution (RS-SISR) techniques have gained significant interest. In this paper, we propose Swin2-MoSE model, an enhanced version of Swin2SR. Our model introduces MoE-SM, an enhanced Mixture-of-Experts (MoE) to replace the Feed-Forward inside all Transformer block. MoE-SM is designed with Smart-Merger, and new layer for merging the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one. Furthermore, we analyze how positional encodings interact with each other, demonstrating that per-channel bias and per-head bias can positively cooperate. Finally, we propose to use a combination of Normalized-Cross-Correlation (NCC) and Structural Similarity Index Measure (SSIM) losses, to avoid typical MSE loss limitations. Experimental results demonstrate that Swin2-MoSE outperforms SOTA by up to dB (PSNR) on task of , and resolution-upscaling ( and OLI2MSI datasets). We show the efficacy of Swin2-MoSE, applying it to a semantic segmentation task (SeasoNet dataset). Code and pretrained are available on https://github.com/IMPLabUniPr/swin2-mose/tree/official_code
1 Introduction
@@
In recent years, the importance of image Super-Resolution (SR), which involves the creation of high-quality images from Low-Resolution (LR) ones, has gained considerable attention from the scientific community [1, 2, 3]. The primary goal of SR is to convert blurred, fuzzy, or indistinct LR images with unrefined details into clear and visually appealing high-resolution (HR) images with enhanced details. This augmentation of image details can be beneficial for several image processing and machine learning tasks, including image classification and segmentation.
@@
Considerable research efforts have been dedicated to advancing super-resolution techniques in the more specific field of remote sensing imagery [4, 5, 6, 7]. The literature firstly focused on rectifying the inherently low spatial resolution in images obtained from UAVs or satellites. Nowadays, with the great availability of satellites that are capable of capturing a multitude of additional spectral bands, such as the near-infrared and infrared spectra, this issue of super-resolution is particularly relevant. While such multi-spectral imagery offers an increased spectral range of frequencies per-pixel, it concurrently introduces the trade-off of further reduced spatial resolution. Indeed, the spatial and spectral resolution of Earth Observation satellites typically do not meet the desired high level due to constraints imposed by optical and sensor technology, as well as the associated high costs. On the other side, many applications could benefit from the availability of higher resolution images, e.g. land use and land cover classification [8], object detection [9], change detection [10], resource management of natural resources as forests, water, and minerals. For this reason, it is very attractive to have methods that allow the reconstruction of missing information through image SR.
@@
SR techniques can be categorized mainly into two types: multi-image super-resolution (MISR) and single-image super-resolution (SISR). MISR involves the reconstruction of a HR image from a collection of LR images, which are typically obtained from different viewpoints or at different times during satellite passes [11, 12, 13]. In this work, the discussion will focus on SISR, a technique in which the HR image is generated starting from a single LR input. Compared to MISR, SISR holds particular appeal for several reasons. Firstly, it entails a reduction in complexity, as it only needs to process a single image and therefore does not require the alignment and registration of multiple images, which can be computationally expensive. Secondly, it can be applied to a wider range of scenarios compared to MISR, as it does not require multiple images of the same scene. Finally, it is easier to implement because it only requires the acquisition of a single image, and less prone to introduce blurring or loss of detail due to the interpolation process. At the same time, while both MISR and SISR are considered as ill-posed problems because there exist an infinite number of possible HR images that could correspond to a given LR image, in the case of SISR, the task is more challenging because, by definition, we only have one image available.
@@
A new End-to-End architecture for Single-Image Super Resolution (SISR) called Swin2-MoSE (Swin V2 Mixture of Super-resolution Experts), improving Swin2SR [14] baseline.
The introduction of MoE-SM, an enhanced Sparsely-Gated Mixture-of-Experts (MoE) layer [15], that replaces the Feed-Forward networks inside Transformer building blocks, with the aim of increasing performance. It is designed with a new advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one.
@@
Evaluation of the proposed model effectiveness for SR imaging in the context of semantic segmentation, utilizing the SeasoNet [8] dataset.
@@ 2 Related Works
Single-Image Super-Resolution. A super-resolution technique tries to reconstruct a high-resolution image from a low-resolution observation. Pioneering the resolution enhancement task, SRCNN [20] employed a three-layer convolutional structure to learn the mapping between LR and HR images. Subsequent advancements, namely RCAN [21] and VDSR [22], addressed the challenges of optimizing deep networks through the incorporation of residual learning techniques. Despite the lightweight nature of these approaches, the convolution kernel usually has a limited receptive field and hinders the extraction of long-range or non-local features. As a result, for some regions with fine details, these methods exhibit sub-optimal performance. With the emergence of the Generative Adversarial Network (GAN) framework, SRGAN [23] was introduced showcasing its ability in super-resolution tasks by generating photo-realistic images, even for large upscaling factors, leveraging content, perceptual, and adversarial losses. ESRGAN [24] refines the latter method by removing Batch Normalization layer, introducing Residual-in-Residual Dense Blocks, and enhancing the perceptual loss through Mean Squared Error (MSE) calculations on high-level feature maps before activation. For capturing long-distance spatial contextual information Second-order Attention Network (SAN) [21] introduces a non-locally enhanced residual group and implements a novel second-order channel attention module to learn feature inter-dependencies by considering second-order statistics of features. Holistic Attention Network (HAN) [25] improves inter-layer dependencies through a Layer Attention Module and augments channel and spatial inter-dependencies using a Channel-Spatial Attention Module, addressing some issues from previous method. The UFormer [26] introduces U-shaped encoder-decoder structure with a window-based system to alleviate self-attention calculation, similar to the Swin Transformer [17] approach. Building upon the Swin transformer architecture, authors of SwinIR [27] introduced a novel Residual Swin Transformer Block for deep feature extraction exhibiting noteworthy performance for this task. In the SwinFIR [28], authors replaced the convolutions () inside the SwinIR with a Fast Fourier Convolution component to have the image-wide receptive field. Innovating upon SwinIR, Swin2SR [14] incorporates the Swin Transformer V2 [29] block, enhancing scalability for higher resolution input images, ensuring stability during training, and improving overall efficiency. Our model extends Swin2SR functionality using a new combination of positional encoding inside the self-attention and introducing an enhanced Sparse Mixture of Experts building block, called MoE-SM, instead of MLP inside the Transformer block. Finally, we also introduced a new loss combination for training our model.
@@
Remote Sensing Single-Image Super-Resolution. In prior investigations, conventional super-resolution techniques like SRCNN [30] and VDSR [31], employing IHS (Intensity, Hue, Saturation) transformations, have demonstrated promising performance when applied to very high resolution satellite data obtained through SPOT6/7 and Pleiades 1A/1B platforms. However, due to the intricate nature of remote sensing imagery, it becomes imperative to design specialized models tailored specifically for Remote Sensing Single-Image Super-Resolution (RS-SISR) [32]. To address this requirement, Lei et al. [33] developed Local-Global Combined Network, featuring a local-global fusion architecture that leverages multiple levels of representation learned via concatenation among various convolutional layers within the network. Xu et al. [34] presented Deep Memory Connected Network, which integrates both image details and environmental context into its framework, utilizing local and global memory connections. Ren et al. [35] introduced Enhanced Residual Convolutional Neural Network with Dual-Luminance Scheme, enhancing the feature flow module’s capability while improving the distinction between features across different map scales. Additionally, Generative Adversarial Networks (GANs) have been employed in real-world RS-SISR applications, as exemplified by Guo et al.’s [36] introduction of Novel Dense GAN alongside their Real High-Resolution and Low-Resolution Aerial Imagery dataset, consisting of paired high-resolution and low-resolution aerial photographs captured under varying elevations. Furthermore, Zhang et al. [37] introduced an integrated framework to simulate remote sensing image degradation, encompassing blur kernel estimation and noise pattern extraction. They utilize a Residual Balanced Attention Network along with a modified U-NET model acting as a discriminator, thereby boosting perceptual performance. Xiao et al. [38] devised a Cross-scale Hierarchical Transformer (CHT) that excels in modeling both global and local cross-scale representations. CHT employs two distinct modules — cross-scale self-attention and cross-scale channel attention — to discern meaningful patterns across varied spatial scales. Tu et al. [39] suggested a Generative Adversarial Network incorporating a fusion of convolutional and Swin Transformer layers for super-resolution reconstruction of remote sensing images. Their method emphasizes large-sized, informative, and correlated pixel properties specific to remote sensing imagery.
@@
Positional Encoding. In Transformer models, since self-attention is invariant to token permutations and does not consider positional information of tokens, positional encodings are commonly utilized to reintroduce this lost positional data. Several approaches have been designed, including Absolute Positional Encoding (APE) [40], Relative Positional Encoding (RPE) [41, 17] and Locally-Enhanced Positional Encoding (LePE) [16]. The first one is also the most used and consists in a sinusoidal function added to the inputs to inject the absolute position of the token. The second encoding (RPE) focuses on capturing the pairwise relationships between elements in the input sequence. This method is particularly useful for tasks where the relative ordering or distance between elements is important and adds the translation equivariance property, similar to convolutions. The last one (LePE) introduces a per-channel bias, rather than the RPE which implement a per-head bias. Similarly to RPE, in Swin Transformer v2 [29, 14], a log-spaced Continuous Position Bias (log-spaced CPB or log CPB) approach is introduced to improve extrapolation accuracy when transferring between large size changes across windows sizes, by compressing large value ranges into a smaller, more manageable scale. For this purpose, they adopted a Continuous Relative Position bias, integrating a small MLP for translating relative coordinates. Since RPE and LePE possess interesting and useful characteristics for the super-resolution task, both were used in our model, obtaining both relative position information and improved locality information.
@@
Mixture-of-Experts (MoE). The Mixture-of-Experts approach was introduced in natural language processing (NLP) within a recurrent language model [15]. It was used to implement conditional computation, allowing the model to scale up to billions of weights, while limiting the increase in computational cost. Sparse-Mixture-of-Experts models [42] are a type of neural network architecture that uses a subset of specialized sub-networks (experts) to improve performance on a given task. The sparsity [15], obtained by training a gate network, permitted to increase the model size significantly, without sacrificing efficiency too much. In a related study [43], authors use a small regression network for image degradation modelling. Simultaneously, multiple convolutional experts are jointly trained within the MoE framework to collectively cover the degradation space. During evaluation, adaptive combination of expert parameters yields a singular optimized network for inference, by that enhancing efficiency. In the paper [44], authors use specialized low-frequency and high-frequency MoE blocks together to ensure the efficient reconstruction of both low-frequency and high-frequency information. In the paper [45], authors introduced MoE inside a Swin Transformer model for the classification task, by replacing the MLP, except for the first two network stage. Similarly to Switch Transformer [46], SwinV2-MoE [45] and V-MoE [47], we decided to replace the MLP of the Transformer block with a MoE. Differently from the first, we made the design choice to use more than one expert each time, with the resulting routing algorithm. While in SwinV2-MoE the first two network stage were not replaced, in our model all the MLP layers are substituted with MoE. Moreover, differently from all of them, we chose a per-example strategy instead of a per-token one to feed the experts, and we designed a new way to merge experts output with a Smart Merger (SM) layer.
@@ 3.1 Overall Network Architecture
Figure 1 shows the overall design of the Swin2-MoSE system. Based on the limitations of the original Swin2SR [14] model, the following subsections will describe the changes proposed in our model to improve the ability to perform super-resolution tasks. In subsection 3.2, we propose the introduction of a new MoE module (called MoE-SM) inside the Transformer blocks (called S2ML, Swin v2 MoSE Layer). Then, in subsection 3.3, how positional encodings interact with each other will be discussed, and some conclusions about their use on Transformer architecture for the super-resolution task will be drawn. Finally, in subsection 3.4, we also investigate various loss functions to improve the robustness and performance of the model training.

### v2

# Leonardo Rossi et al

@@ Abstract
[Abstract] Due to the limitations of current optical and sensor technologies and the high cost of updating them, the spectral and spatial resolution of satellites may not always meet desired requirements. For these reasons, Remote-Sensing Single-Image Super-Resolution (RS-SISR) techniques have gained significant interest. In this paper, we propose Swin2-MoSE model, an enhanced version of Swin2SR. Our model introduces MoE-SM, an enhanced Mixture-of-Experts (MoE) to replace the Feed-Forward inside all Transformer block. MoE-SM is designed with Smart-Merger, and new layer for merging the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one. Furthermore, we analyze how positional encodings interact with each other, demonstrating that per-channel bias and per-head bias can positively cooperate. Finally, we propose to use a combination of Normalized-Cross-Correlation (NCC) and Structural Similarity Index Measure (SSIM) losses, to avoid typical MSE loss limitations. Experimental results demonstrate that Swin2-MoSE outperforms any Swin derived models by up to dB (PSNR) on task of , and resolution-upscaling ( and OLI2MSI datasets). It also outperforms SOTA models by a good margin, proving to be competitive and with excellent potential, especially for complex tasks. Additionally, an analysis of computational costs is also performed. Finally, we show the efficacy of Swin2-MoSE, applying it to a semantic segmentation task (SeasoNet dataset). Code and pretrained are available on https://github.com/IMPLabUniPr/swin2-mose/tree/official_code
[…]
keywords:
Single-Image Super-Resolution; Multi-spectral; Remote-Sensing; Earth Observation; Swin Transformer; Mixture-of-Experts; Positional Embedding1 Introduction
Remote sensing has become more and more important recently for its several applications, especially in the agricultural field. Crop segmentation and classification, disease detection, automatic irrigation control are some of the possible applications of multi-spectral satellite images that can increase crops productivity and help to fight the waste of resources. Nevertheless, due to optical and sensor technology, satellites often provide images whose resolution is not enough for their target applications. For this reason, studying new and powerful remote sensing super resolution techniques is of paramount importance and with this work we were able to provide a novel method that outperforms the current state of the art.
@@
In recent years, the importance of image Super-Resolution (SR), which involves the creation of high-quality images from Low-Resolution (LR) ones, has gained considerable attention from the scientific community 1, 2, 3. The primary goal of SR is to convert blurred, fuzzy, or indistinct LR images with unrefined details into clear and visually appealing high-resolution (HR) images with enhanced details. This augmentation of image details can be beneficial for several image processing and machine learning tasks, including image classification and segmentation.
@@
Considerable research efforts have been dedicated to advancing super-resolution techniques in the more specific field of remote sensing imagery 4, 5, 6, 7. The literature firstly focused on rectifying the inherently low spatial resolution in images obtained from UAVs or satellites. Nowadays, with the great availability of satellites that are capable of capturing a multitude of additional spectral bands, such as the near-infrared and infrared spectra, this issue of super-resolution is particularly relevant. While such multi-spectral imagery offers an increased spectral range of frequencies per-pixel, it concurrently introduces the trade-off of further reduced spatial resolution. Indeed, the spatial and spectral resolution of Earth Observation satellites typically do not meet the desired high level due to constraints imposed by optical and sensor technology, as well as the associated high costs. On the other side, many applications could benefit from the availability of higher resolution images, e.g. land use and land cover classification 8, object detection 9, change detection 10, resource management of natural resources as forests, water, and minerals. For this reason, we decided to focus on the development of a framework which addresses the lack of high resolution multispectral imagery in the context of remote sensing, combining the large availability of low resolution data with a method that allows to reconstruct reliably high frequency details with respect to previous methods. Moreover, addressing the scarcity of super-resolution methods in remote sensing area, we tailor our framework specifically for satellite imagery which show very different scale and semantics compared to regular RGB imagery on which the vast majority of SR methods are trained for.
@@
SR techniques can be categorized mainly into two types: multi-image super-resolution (MISR) and single-image super-resolution (SISR). MISR involves the reconstruction of a HR image from a collection of LR images, which are typically obtained from different viewpoints or at different times during satellite passes 11, 12, 13. In this work, the discussion will focus on SISR, a technique in which the HR image is generated starting from a single LR input. Compared to MISR, SISR holds particular appeal for several reasons. Firstly, it entails a reduction in complexity, as it only needs to process a single image and therefore does not require the alignment and registration of multiple images, which can be computationally expensive. Secondly, it can be applied to a wider range of scenarios compared to MISR, as it does not require multiple images of the same scene. Finally, it is easier to implement because it only requires the acquisition of a single image, and less prone to introduce blurring or loss of detail due to the interpolation process. At the same time, while both MISR and SISR are considered as ill-posed problems because there exist an infinite number of possible HR images that could correspond to a given LR image, in the case of SISR, the task is more challenging because, by definition, we only have one image available.
@@
A new End-to-End architecture for Single-Image Super Resolution (SISR) called Swin2-MoSE (Swin V2 Mixture of Super-resolution Experts), improving Swin2SR 14 baseline.
The introduction of MoE-SM, an enhanced Sparsely-Gated Mixture-of-Experts (MoE) layer 15, that replaces the Feed-Forward networks inside Transformer building blocks, with the aim of increasing performance. It is designed with a new advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one.
@@ 2 Related Works
Single-Image Super-Resolution. A super-resolution technique tries to reconstruct a high-resolution image from a low-resolution observation. Pioneering the resolution enhancement task, SRCNN 18 employed a three-layer convolutional structure to learn the mapping between LR and HR images. Subsequent advancements, namely RCAN 19 and VDSR 20, addressed the challenges of optimizing deep networks through the incorporation of residual learning techniques. Despite the lightweight nature of these approaches, the convolution kernel usually has a limited receptive field and hinders the extraction of long-range or non-local features. As a result, for some regions with fine details, these methods exhibit sub-optimal performance. With the emergence of the Generative Adversarial Network (GAN) framework, SRGAN 21 was introduced showcasing its ability in super-resolution tasks by generating photo-realistic images, even for large upscaling factors, leveraging content, perceptual, and adversarial losses. ESRGAN 22 refines the latter method by removing Batch Normalization layer, introducing Residual-in-Residual Dense Blocks, and enhancing the perceptual loss through Mean Squared Error (MSE) calculations on high-level feature maps before activation. For capturing long-distance spatial contextual information Second-order Attention Network (SAN) 19 introduces a non-locally enhanced residual group and implements a novel second-order channel attention module to learn feature inter-dependencies by considering second-order statistics of features. Holistic Attention Network (HAN) 23 improves inter-layer dependencies through a Layer Attention Module and augments channel and spatial inter-dependencies using a Channel-Spatial Attention Module, addressing some issues from previous method. The UFormer 24 introduces U-shaped encoder-decoder structure with a window-based system to alleviate self-attention calculation, similar to the Swin Transformer 17 approach. Building upon the Swin transformer architecture, authors of SwinIR 25 introduced a novel Residual Swin Transformer Block for deep feature extraction exhibiting noteworthy performance for this task. In the SwinFIR 26, authors replaced the convolutions () inside the SwinIR with a Fast Fourier Convolution component to have the image-wide receptive field. Innovating upon SwinIR, Swin2SR 14 incorporates the Swin Transformer V2 27 block, enhancing scalability for higher resolution input images, ensuring stability during training, and improving overall efficiency. All these models, descending from a Swin architecture, contain a small MLP within the Transformer block which may not be able to adapt effectively to the growth of available data. Inspired by the successes of MoE 28, 29, 15 architectures, we have successfully explored its use in super-resolution, introducing an enhanced Sparse Mixture of Experts building block, called MoE-SM, instead of MLP inside the Transformer block. Authors of DAT 30 proposed to use spatial and channel self-attention together to exploit both context information. Like them, we too felt the need for spatial and channel information. In our case however, we use two different positional encodings within the same self-attention. In addition, unlike the others, in our work we explore the use of a combination of losses, to overcome the limits related to overfit on low frequencies of the signal.
@@
Remote Sensing Single-Image Super-Resolution. In prior investigations, conventional super-resolution techniques like SRCNN 31 and VDSR 32, employing IHS (Intensity, Hue, Saturation) transformations, have demonstrated promising performance when applied to very high resolution satellite data obtained through SPOT6/7 and Pleiades 1A/1B platforms. However, due to the intricate nature of remote sensing imagery, it becomes imperative to design specialized models tailored specifically for Remote Sensing Single-Image Super-Resolution (RS-SISR) 33. To address this requirement, Lei et al. 34 developed Local-Global Combined Network, featuring a local-global fusion architecture that leverages multiple levels of representation learned via concatenation among various convolutional layers within the network. Xu et al. 35 presented Deep Memory Connected Network, which integrates both image details and environmental context into its framework, utilizing local and global memory connections. Ren et al. 36 introduced Enhanced Residual Convolutional Neural Network with Dual-Luminance Scheme, enhancing the feature flow module’s capability while improving the distinction between features across different map scales. Additionally, Generative Adversarial Networks (GANs) have been employed in real-world RS-SISR applications, as exemplified by Guo et al.’s 37 introduction of Novel Dense GAN alongside their Real High-Resolution and Low-Resolution Aerial Imagery dataset, consisting of paired high-resolution and low-resolution aerial photographs captured under varying elevations. Furthermore, Zhang et al. 38 introduced an integrated framework to simulate remote sensing image degradation, encompassing blur kernel estimation and noise pattern extraction. They utilize a Residual Balanced Attention Network along with a modified U-NET model acting as a discriminator, thereby boosting perceptual performance. Xiao et al. 39 devised a Cross-scale Hierarchical Transformer (CHT) that excels in modeling both global and local cross-scale representations. CHT employs two distinct modules — cross-scale self-attention and cross-scale channel attention — to discern meaningful patterns across varied spatial scales. Tu et al. 40 suggested a Generative Adversarial Network incorporating a fusion of convolutional and Swin Transformer layers for super-resolution reconstruction of remote sensing images. Their method emphasizes large-sized, informative, and correlated pixel properties specific to remote sensing imagery.
@@
Positional Encoding. In Transformer models, since self-attention is invariant to token permutations and does not consider positional information of tokens, positional encodings are commonly utilized to reintroduce this lost positional data. Several approaches have been designed, including Absolute Positional Encoding (APE) 41, Relative Positional Encoding (RPE) 42, 17 and Locally-Enhanced Positional Encoding (LePE) 16. The first one is also the most used and consists in a sinusoidal function added to the inputs to inject the absolute position of the token. The second encoding (RPE) focuses on capturing the pairwise relationships between elements in the input sequence. This method is particularly useful for tasks where the relative ordering or distance between elements is important and adds the translation equivariance property, similar to convolutions. The last one (LePE) introduces a per-channel bias, rather than the RPE which implement a per-head bias. Similarly to RPE, in Swin Transformer v2 27, 14, a log-spaced Continuous Position Bias (log-spaced CPB or log CPB) approach is introduced to improve extrapolation accuracy when transferring between large size changes across windows sizes, by compressing large value ranges into a smaller, more manageable scale. For this purpose, they adopted a Continuous Relative Position bias, integrating a small MLP for translating relative coordinates. Since RPE and LePE possess interesting and useful characteristics for the super-resolution task, unlike what others do, both were used in our model, obtaining both relative position information and improved locality information.
@@
Mixture-of-Experts (MoE). The Mixture-of-Experts approach was introduced in natural language processing (NLP) within a recurrent language model 15. It was used to implement conditional computation, allowing the model to scale up to billions of weights, while limiting the increase in computational cost. Sparse-Mixture-of-Experts models 29 are a type of neural network architecture that uses a subset of specialized sub-networks (experts) to improve performance on a given task. The sparsity 15, obtained by training a gate network, permitted to increase the model size significantly, without sacrificing efficiency too much. In a related study 43, authors use a small regression network for image degradation modelling. Simultaneously, multiple convolutional experts are jointly trained within the MoE framework to collectively cover the degradation space. During evaluation, adaptive combination of expert parameters yields a singular optimized network for inference, by that enhancing efficiency. In the paper 44, authors use specialized low-frequency and high-frequency MoE blocks together to ensure the efficient reconstruction of both low-frequency and high-frequency information. In the paper 45, authors introduced MoE inside a Swin Transformer model for the classification task, by replacing the MLP, except for the first two network stage. Similarly to Switch Transformer 46, SwinV2-MoE 45 and V-MoE 28, we decided to replace the MLP of the Transformer block with a MoE. Differently from the first, we made the design choice to use more than one expert each time, with the resulting routing algorithm. While in SwinV2-MoE the first two network stages were not replaced, in our model all the MLP layers are substituted with MoE. Due to spatial proximity in images, it is very likely that nearby tokens are similar, in our opinion it is much more favorable that all the tokens of an example are processed by the same experts. or this reason, differently from the others, we have chosen a per-example strategy instead of a per-token one to feed the experts. This should facilitate the specialization ability of each expert and reduce the fragmentation of experts in the case of distributed training. Furthermore, one of the aspects less explored by other methods is linked to the fusion of the results of individual experts. For this reason, we designed a new way to merge experts output with a Smart Merger (SM) layer.
@@ 3.1 Overall Network Architecture
Figure 1 shows the overall design of the Swin2-MoSE system. Based on the limitations of the original Swin2SR 14 model, the following subsections will describe the changes proposed in our model to improve the ability to perform super-resolution tasks. In subsection 3.2, we propose the introduction of a new MoE module (called MoE-SM) inside the Transformer blocks (called S2ML, Swin v2 MoSE Layer). Then, in subsection 3.3, how positional encodings interact with each other will be discussed, and some conclusions about their use on Transformer architecture for the super-resolution task will be drawn. Finally, in subsection 3.4, we also investigate various loss functions to improve the robustness and performance of the model training.